OpenStudy (anonymous):

Identify the slope and y-intercept of the inequality 8x + 2y > 12. slope = 4, y-intercept = 6 slope = -4, y-intercept = 12 slope = -4, y-intercept = 6 slope = -1/4, y-intercept = 12

jimthompson5910 (jim_thompson5910):

hint: solve 8x + 2y = 12 for y tell me what you get when you do so

OpenStudy (anonymous):

what value do i make y to do that

jimthompson5910 (jim_thompson5910):

what do you mean?

OpenStudy (anonymous):

i dont know how to do that can you how me please :))))

jimthompson5910 (jim_thompson5910):

sure I'll show you how to do it so you can apply it to similar problems in the future

jimthompson5910 (jim_thompson5910):

the goal is to isolate y, ie get it by itself you do this by subtracting 8x from both sides, then by dividing both sides by 2 like this 8x + 2y = 12 8x + 2y - 8x = 12 -8x 2y = 12 -8x 2y = -8x + 12 2y/2 = (-8x + 12)/2 y = -8x/2 + 12/2 y = -4x + 6

jimthompson5910 (jim_thompson5910):

so we go from 8x + 2y = 12 to y = -4x + 6

jimthompson5910 (jim_thompson5910):

from here, you should be able to see the slope and y-intercept

jimthompson5910 (jim_thompson5910):

does that all make sense?

OpenStudy (anonymous):

sorry i didnt see your comment yes i get that step so do you add 6 on both sides now

jimthompson5910 (jim_thompson5910):

no y = -4x + 6 is in slope intercept form

jimthompson5910 (jim_thompson5910):

slope intercept form is y = mx+b

jimthompson5910 (jim_thompson5910):

so what is the slope m?

OpenStudy (anonymous):

mtae... y=mx+b where m is slope and b is y intercept

OpenStudy (anonymous):

What is the slope of y = -4x + 6?